LBC+LUCIFER Technical Observing, prep for Partner Science: UT 2010 January 31

Observer: DThompson (LBTO), OKuhn, ARakich (Tucson)
Telescope Operator: WWack
Instrument Support: JMorris, NAgeorges
Telescope Support: JUrban
SW support: TLeibold (on call)


  • Observed ~1.5 hours with LBC
  • Lost remainder of night to clouds, fog, high humidity



Heavy clouds at sunset.

UT04:22 Still cloudy. RH=49% (dewpoint is rapidly rising starting 30 minutes ago). Toutside=-4.4, Tmirror=-1.8. Vent doors open and mirror ventilation on since sunset.

UT06:05 Dome open, left side only (right shutter did not work). Still fairly cloudy, but some stars visible. Olga and Andrew called.

UT7:50 Dome closed, rapidly rising humidity. Olga and Andrew did some more testing of DOFPIA but have given up for the night. When the dome is closed Wayne will reconfigure for LUCIFER in case it clears back up.

UT12:30 Still foggy.
UT12:45 Calling it a night.

Observing Plan

  • Finish up LBC FPIA tests from previous night, then switch to LUCIFER

  • Correct initial collimation/pointing with IE/CA (and IA?)
    • modify/test ptcorr script to use pcsutilities command
    • example: pcsutilities ptmodify ca -200.33

  • Check existing collimation model

  • Check existing pointing model

  • Run short dither (dice5) pattern to check that things still work

  • Re-test updatePointingReference functionality.

  • Collect through-focus sequence for Andrew - reasonably dense star field.

  • Collect astrometric field data to check AGw calibration at colder temps

  • Practice MOS acquisitions with Nancy
    • Read appropriate section in documentation

  • Execute MOS preimaging for partnership. Scripts received from:
    • Ian McGreer

-- DavidThompson - 30 Jan 2010
Topic revision: r6 - 31 Jan 2010, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback